# Quantitative surgery and total mean curvature

### v1

@@ 2. Total mean curvature bounds with fixed positive lower mean curvature bound
Figure 6. The end of the local deformation $\Psi(1)$ obtained by applying the local flexibility lemma to the deformation $\psi$.
We set $g_{\widetilde{\Omega}_{i}}\coloneqq\Psi(1)$. Observe, that the metric $g_{\widetilde{\Omega}_{i}}$ on $\widetilde{\Omega}_{i}\cap\Omega=\Omega$, the metric $h^{\prime}$ on $\widetilde{\Omega}_{i}\cap(U_{i}\times(-\delta^{\prime},1))$ and the metric $h$ on $\widetilde{\Omega}_{i}\cap{\rm int}(W)$ are compatible with each other. Hence, the union of these metrics defines a smooth metric $g_{\widetilde{\Omega}_{i}}$ on $\widetilde{\Omega}_{i}$. By construction, $H_{g_{\widetilde{\Omega}_{i}}}=H_{g_{i}}=H_{g_{\Omega}}$ on $\partial\Omega\setminus Z_{i}$. Furthermore, we have $\mathrm{scal}_{g_{\widetilde{\Omega}_{i}}}>{\rm min}\{\sigma,\mathrm{scal}_{h}\}-1$ for $i\in\{0,1\}$. Thus, $g_{\widetilde{\Omega}_{i}}$ satisfies the properties required above.
3. Uniform upper bound of total mean curvature for families of metrics
@@
Let $(M,g_{M})$ be a closed spin Riemannian manifold for which Theorem 1.2 holds. In this section we will prove that Theorem 1.2 still holds, with a uniform upper bound $\Lambda$, for a certain family of metrics on $M$. The proof is based on arguments in [25].
@@
Let $s\in\mathbb{R}$ such that $\mathrm{scal}_{g}>s$. We define $\mathscr{R}_{g_{M},s}(M)$ as the subset of Riemannian metrics $g$ on $M$ with the following property: There exists a smooth path $\lambda_{g}\colon[0,1]\to C^{\infty}({\rm Sym}(T^{*}M\otimes T^{*}M))$ of Riemannian metrics starting at $g$ and ending at $g_{M}$ such that, for all $t\in[0,1]$, we have
@@
    $\lambda^{\prime}_{g}(t)>0$ in the sense of symmetric $(0,2)$-tensor fields.
@@
    $\mathrm{scal}_{\lambda_{g}(t)}>s$ on $M$.
@@
For any $s\in\mathbb{R}$, there exists $\Lambda=\Lambda(M,g_{M},\sigma,s)\in\mathbb{R}_{+}$ such that the following holds: If $g\in\mathscr{R}_{g_{M},s}$ and if $(\Omega,g_{\Omega})$ is a fill-in of $(M,g)$ with $\mathrm{scal}_{g_{\Omega}}\geq\sigma$ and $H_{g_{\Omega}}\geq 0$, then
$\int_{\partial\Omega}H_{g_{\Omega}}{\rm dvol}_{g_{M}}\leq\Lambda(M,g_{M},\sigma,s).$
@@
Let $g\in\mathscr{R}_{g_{M},s}$. Let $\lambda_{g}\colon[0,1]\to C^{\infty}({\rm Sym}(T^{*}M\otimes T^{*}M))$ be a smooth path from $g$ to $g_{M}$ that satisfies properties (i) and (ii). Consider the smooth metric $\overline{g}$ on $M\times[0,1]$ defined by
$\overline{g}=dt^{2}+\lambda_{g}(t).$
Let $\overline{A}_{t}$ be the second fundamental form and $\overline{H}_{t}=\operatorname{tr}_{\lambda_{g}(t)}\overline{A}_{t}$ be the mean curvature of $\Sigma_{t}:=M\times\{t\}\subset(M\times[0,1],\overline{g})$ with respect to the unit normal field $\frac{\partial}{\partial t}$. By (i), we have $\overline{A}_{t}>0$, hence $\overline{H}_{t}>0$ and $\overline{H}_{t}^{2}-\|\overline{A}_{t}\|^{2}>0$ for $t\in[0,1]$
Let $(\Omega,g_{\Omega})$ be a fill-in of $(M,g)$ with $\mathrm{scal}_{g_{\Omega}}\geq\sigma$ and $H_{g_{\Omega}}\geq 0$. Applying [4, Proposition 3.8], we can increase the mean curvature $H_{g_{\Omega}}$ slightly, while replacing $\sigma$ by $\sigma-1$. Hence, we can asume that $H_{g_{\Omega}}>0$. On $M\times[0,1]$, consider the Bartnik-type parabolic PDE
$\overline{H}_{t}\frac{\partial u}{\partial t}$ $=u^{2}\Delta_{\lambda_{g}(t)}u+\frac{1}{2}(s-\mathrm{scal}_{\lambda_{g}(t)})u^{3}+\frac{1}{2}(\mathrm{scal}_{\lambda_{g}(t)}-\mathrm{scal}_{\overline{g}})u$
@@
Since, by (ii), we have $s-\mathrm{scal}_{\lambda_{g}(t)}<0$ for $t\in[0,1]$, this equation has a positive solution $u\colon M\times[0,1]\to\mathbb{R}_{>0}$. Consider the Riemannian metric $\widetilde{g}=u^{2}dt^{2}+\lambda_{g}(t)$ on $M\times[0,1]$.
Let $H_{t}=u^{-1}\overline{H}_{t}>0$ be the mean curvature on $\Sigma_{t}$ induced by $\widetilde{g}$ with respect to the unit normal field $u^{-1}\frac{\partial}{\partial t}$. Note that $H_{0}=H_{g_{\Omega}}$. By the computation in the proof of [25, Theorem 4.1] and using that $\overline{H}_{t}^{2}-\|\overline{A}_{t}\|^{2}>0$, we have
$\frac{\partial}{\partial t}\int_{\Sigma_{t}}H_{t}\,\rm{dvol}_{\lambda_{g}(t)}>0.$
@@
(3.1) $\int_{\Sigma_{1}}H_{1}{\rm dvol}_{\lambda_{g}(1)}>\int_{\Sigma_{0}}H_{0}{\rm dvol}_{\lambda_{g}(0)}=\int_{M}H_{g_{\Omega}}{\rm dvol}_{g}.$
Furthermore, we have $\mathrm{scal}_{\widetilde{g}}={\rm const.}=s$.
Consider the topological manifold
$\widetilde{\Omega}:=\Omega\bigcup_{M\times\{0\}\approx\partial\Omega}M\times[0,1].$
It carries a unique smooth structure induced by the smooth structures on $\Omega$ and $M\times[0,1]$ and the collar neighborhoods along $\partial\Omega\subset\Omega$ and $M\times\{0\}\subset M\times[0,1]$ induced by $g_{\Omega}$ and $\widetilde{g}$. If $n\geq 7$, the manifold $\widetilde{\Omega}$ carries a spin structure restricting to the given spin structure on $\Omega$. By [4, Theorem 4.11] and by the choice of $u(\cdot,0)$, the mean convex singularity at the gluing region $\partial\Omega=M\times\{0\}$ can be smoothened, that is, there exists a smooth Riemannian metric $g_{\widetilde{\Omega}}$ on $\widetilde{\Omega}$ which is equal to $\widetilde{g}$ near $M\times\{1\}$ and whose scalar curvature is bounded below by $\min\{s,\sigma\}-1$.
@@
$\int_{M}H_{g_{\Omega}}{\rm dvol}_{g}<\int_{\Sigma_{1}}H_{1}{\rm dvol}_{\lambda_{g}(1)}=\int_{\partial\widetilde{\Omega}}H_{g_{\widetilde{\Omega}}}{\rm dvol}_{g_{M}}.$
Now the claim of Proposition 3.2 follows from the assumption on $(M,g_{M})$, with $\sigma$ replaced by ${\rm min}\{\sigma,s\}-1$. ∎
4. Surgery bending in Euclidean space
In this section we construct hypersurfaces $\overline{\Sigma}_{\varrho}$ in Euclidean space which we will later use to define surgery bendings with appropriate curvatures. For this purpose we equip $B_{1}^{n-k}(0)\times[0,1]\subset\mathbb{R}^{\,n-k+1}$ with the flat metric $g_{\mathbb{R}^{n-k+1}}=\mathrm{d}x_{1}^{2}+\dots+\mathrm{d}x_{n-k}^{2}+\mathrm{d}t^{2}$. Let us now describe the construction of $\overline{\Sigma}_{\varrho}$.
Construction 4.1.
Let $\varrho\in(0,\frac{1}{10}]$ be fixed and consider the function
from Lemma A.1. The handle $\overline{\Sigma}_{\varrho}$ will essentially be given as the graph of the function $(x_{1},\dots,x_{n-k})\mapsto\xi_{\varrho}(\sqrt{x_{1}^{2}+\dots+x_{n-k}^{2}})+a_{\varrho}$, where $a_{\varrho}\coloneqq-\xi_{\varrho}(2\sqrt{\varrho})$. However, there are a few adjustments to make: Since we want to glue this hypersurface into a given manifold, we need $\overline{\Sigma}_{\varrho}$ to agree with $B_{1}^{n-k}(0)\times\{0\}$ for large radii, and we need it to be cylindrical near the top.
In order to achieve this, we fix once and for all a cutoff-function $\eta\colon\mathbb{R}\to[0,1]$, that is $\eta\equiv 1$ on $(-\infty,0]$, $\eta\equiv 0$ on $1,\infty)$, and $2\leq\eta^{\prime}\leq 0$, $|\eta^{\prime\prime}|\leq 10$. We now define (see Fig. 7) $\eta_{\varrho}(r)=\eta(\tfrac{r-\sqrt{\varrho}}{\sqrt{\varrho}})$ and
and we consider the graph of the function
$B_{1}^{n-k}(0)\setminus B_{\varrho^{4}}^{n-k}(0)\to\mathbb{R},\qquad(x_{1},\dots,x_{n-k})\mapsto f_{\varrho}\left(\sqrt{x_{1}^{2}+\dots+x_{n-k}^{2}}\right).$
@@ 4. Surgery bending in Euclidean space
Figure 7. The function ${f}_{\varrho}$.
Note that the derivative of $f_{\varrho}$ converges to -$\infty$ for $r\to\varrho^{4}$ and therefore, we can attach a straight cylinder of radius $\varrho^{4}$ to this graph at the level $t=a_{\varrho}=f(\varrho^{4})$ and obtain a $C^{1}$-hypersurface which is smooth away from $t=a_{\varrho}$.
[…]
Figure 8. The hypersurface $\overline{\Sigma}_{\varrho}$ as the graph of a spherically symmetric function.
@@
Let us motivate the choice of the function $\xi_{\varrho}$. The graph of the functions $\ell_{\varrho}(r)=-2\sqrt{2\varrho}\sqrt{r-2\varrho}$ give rise to spherically symmetric surfaces satisfying
$2\lambda_{1}=\lambda_{2}=\dots=\lambda_{n-k}$
for their principal curvatures, cf. Lemma A.3. In $n=3$ these are the (spatial) Schwarzschild metrics of mass $m=\varrho$. In this case the scalar curvature vanishes which does not allow us absorb the error terms appearing in the handle construction. Therefore, we solve the ODE
$4\lambda_{1}=\lambda_{2}=\dots=\lambda_{n-k}$
@@
Let $\overline{\nu}_{\varrho}$ be the unit normal vector field of $\overline{\Sigma}_{\varrho}$ and denote by $\overline{\nu}_{\varrho}^{\intercal}$ the projection of $\overline{\nu}_{\varrho}$ onto the tangent space of the slices $B_{1}^{n-k}(0)\times\{t\}$, i.e.,
@@
The hypersurfaces $\overline{\Sigma}_{\varrho}$ from 4.1 satisfy the following properties.
@@
    $\overline{\Sigma}_{\varrho}$ is $\mathrm{O}(n-k)$-invariant and satisfies
    $\overline{\Sigma}_{\varrho}\cap(B_{1}^{n-k}(0)\times[2\varrho,1])$ $=\partial B_{\varrho^{4}}^{n-k}(0)\times[2\varrho,1],$
    $\overline{\Sigma}_{\varrho}\cap\Bigl(\bigl(B_{1}^{n-k}(0)\setminus B_{2\sqrt{\varrho}}^{n-k}(0)\bigr)\times[0,1]\Bigr)$ $=\bigl(B_{1}^{n-k}(0)\setminus B_{2\sqrt{\varrho}}^{n-k}(0)\bigr)\times\{0\}.$
@@
    For $r\in[\varrho^{4},\sqrt{\varrho}]$ and away from $t=a_{\varrho}$ we have
    $H(\overline{\Sigma}_{\varrho})>\frac{1}{2}|A(\overline{\Sigma}_{\varrho})|\quad\text{if }n-k\geq 2$
    $\mathrm{scal}(\overline{\Sigma}_{\varrho})>\frac{1}{3}|A(\overline{\Sigma}_{\varrho})|^{2}\quad\text{if }n-k\geq 3.$
@@
    For $r\in[\varrho^{4},\varrho^{2}]$, $n-k\geq 3$ and away from $t=a_{\varrho}$, we have
@@
For the final two properties let $\overline{C}\coloneqq 18(n-k)+{432}$.

### v2

@@ 3. Spin fill-ins and creased initial data sets
$\mathcal{H}_{hyp}(V)=4\int_{\Omega}\Big(\big|\nabla_{i}\psi+\frac{1}{2}k_{ij}e_{j}e_{0}\psi\big|^{2}+\frac{1}{2}\big\langle\psi,(\mu+Je_{0})\psi\big\rangle\Big)\geq 0,$
where $V=|\psi^{\infty}|^{2}_{b}$ and the hyperboloidal mass functional $\mathcal{H}_{hyp}(V)$ is defined in [HJZ25, p. 307].
By appropriate rescalings ($g_{ij}\to\lambda^{2}g_{ij}$, $k_{ij}\to\lambda k_{ij}$) an analogous result holds for initial data sets $(\Omega,g,k)$ asymptotic to $\mathbb{H}^{n+1}_{\sigma_{2}}$ instead of the standard hyperbolic space $\mathbb{H}^{n+1}=\mathbb{H}^{n+1}_{n(n+1)}$, cf. [SWW22, p. 235]. See [CJŁ04, CH04] and [HJZ25, Appendix A] for a detailed discussion and for further references on the hyperboloidal PMT. We emphasize that this result should not be confused with the asymptotically AdS PMT, in which the metric $g$ also asymptotes to the hyperbolic metric but $k$ asymptotes to zero [CMT06, HZ25].
In our situation, the initial data set has significantly more restrictive asymptotics: we have $k=c_{2}g$ at infinity, and the Bartnik extension metric has a very specific form. As pointed out in [CH04, p. 232 and Section 3] as well as [BW12, Appendix A], under these so-called Wang asymptotics the mass
(3.4) $m=\mathcal{H}_{hyp}(V_{0}),\qquad V_{0}=V_{0}(\psi_{0}^{\infty})$
reduces to the mass defined in [Wan01], for a certain choice of Killing spinor $\psi^{\infty}_{0}$. In view of (3.3), this in particular implies $m(r_{0})\geq 0$, which yields the desired total mean curvature bound.
Finally, we explain why the corners at $M$ and $N$ do not pose difficulties. The type of corner arising at $N$ is well understood; see, for instance, [BQ08]. We therefore focus on the corner along $M$. Strictly speaking, the results of [Lin23, KKL25] are formulated for asymptotically flat spin manifolds. However, their arguments are entirely local, and it is straightforward to see that they extend to the asymptotically hyperboloidal spin setting as well. More precisely, given a spinor $\psi$ on $N$, let $\psi_{\pm}$ denote its restrictions to $\Omega_{+}:=\Omega_{1}\cup_{N}\Omega_{2}$ and $\Omega_{-}:=\Omega_{0}$, respectively. Recall Witten’s standard divergence identity,
[…]
Integrating this identity by parts on both $\Omega_{-}$ and $\Omega_{+}$, and then imposing the transmission boundary condition [KKL25, Equation (1.9)]
(3.5) $\psi^{-}=\cosh\frac{\vartheta}{2}\,\psi^{+}+\sinh\frac{\vartheta}{2}\,\nu^{+}e_{0}^{+}\psi^{+},$
[…]
We note that the hyperboloidal positive mass theorem can in fact be reduced to the asymptotically flat case, as shown in [CD19]. Thus, the results of [KKL25] can also be applied directly. This finishes the proof of Theorem (C).
The hyperboloidal positive mass theorem also holds for non-spin manifolds, see [ACG08, Sak21, Lun23, CD19a]. However, the non-smooth case treated in [KKL25] does not yet have an analogue in the non-spin setting. Resolving this issue would extend the statement of Theorem (C) to the non-spin case.
We close this section with some remarks on the asymptotics of $\Lambda(M,g_{M},\sigma,\kappa)$ for fixed $(M,g_{M})$. In [Bär26] it is shown that $\Lambda(M,g_{M},\sigma,\kappa)$ can be chosen to depend linearly on $\sqrt{|\sigma|}$ and $|\kappa|$.
Our constant $\Lambda(M,g_{M},\sigma_{0},\kappa)$, in case $M\approx S^{n}$, can be explicitly estimated via
(3.6) $H(N=S^{n}_{r_{0}}\subseteq\mathbb{H}^{n+1}_{\sigma_{2}})=n\sqrt{-\frac{\sigma_{2}}{n(n+1)}}\coth\left(r_{0}\sqrt{-\frac{\sigma_{2}}{n(n+1)}}\right),$
and
(3.7) $\sigma_{2}=-\frac{n+1}{n}\cosh(1)^{2}\max\left\{16\kappa^{2},\frac{n}{n+1}|\sigma_{0}|,(\cosh 1)^{-1}\frac{n}{n+1}|\sigma_{1}|\right\}.$
For $\sigma_{0}\gg 1$ or $\kappa\gg 1$, we therefore have $H(N\subseteq\mathbb{H}^{n+1}_{\sigma_{2}})\approx\max\{\sqrt{-\sigma_{0}},-\kappa\}$, which gives the same asymptotics as Bär’s proof.
With the methods developed in the succeeding sections, we can show that the same asymptotics hold for all closed spin manifolds. To justify this, let $(\Omega_{0},g_{0})$ be a spin fill-in of $(M,g_{M})$ such that $\mathrm{scal}_{g_{0}}\geq\sigma_{0}$ and $H_{g_{0}}\geq\kappa$ for some $\sigma_{0},\kappa<0$. Let $h_{0}\colon M\to\mathbb{R}_{>0}$ be the (smoothened) function in (3.1). According to the proof of [SWW22, Lemma 3.1 and Theorem 4.1], there exists a Riemannian metric $g_{1}$ on $M\times[0,1]$ such that the cobordism $(\Omega_{1}=M\times[0,1],g_{1})$ connects $(M,g_{M})$ to $(N:=M,2g_{M})$ satisfying properties (i), (ii) and (iii) above. This results in a spin fill-in $(\Omega=\Omega_{0}\cup_{M}\Omega_{1},g_{\Omega})$ with positive mean curvature along $(N,2g_{M})$, with a crease at $M$ and with scalar curvature equal to $\sigma_{1}=\sigma_{1}(M,g_{M})$.
[…]
  • $g_{\widetilde{\Omega}}=g_{\Omega}=g_{0}$ on $\Omega_{0}$,
  • $\mathrm{scal}(\widetilde{\Omega},g_{\widetilde{\Omega}})\geq\min\{\sigma_{0},C-\ell\}$,
[…]
By (3.2) and (3.3), we thus obtain
$\int_{N}H\big(M\subseteq\Omega_{0})\leq 2^{\ell}\cdot 3\cdot\int_{S^{n}_{r_{0}}}H(S^{n}_{r_{0}}\subseteq\mathbb{H}^{n+1}_{\sigma_{2}})$
for some $r_{0}=r_{0}(M,g_{M})>0$ and for $\sigma_{2}$ given by
$\sigma_{2}=-\frac{n+1}{n}\cosh(1)^{2}\max\left\{16\kappa^{2},-\sigma_{0}\frac{n}{n+1},-C+\ell\right\},$
similar to (3.7). Hence, there exists a $C_{0}(M,g_{M})>0$, independent of $\kappa$ and $\sigma_{0}$, (3.6) gives
(3.8) $\Lambda(M,g_{M},\sigma_{0},\kappa)\leq C_{0}(M,g_{M})\cdot\big(\max\{\sqrt{-\sigma_{0}},-\kappa\}+1\big).$
4. Fill-ins of families of metrics
@@
Let $(M,g_{M})$ be a closed oriented Riemannian manifold. In this section we derive a uniform upper bound for the total mean curvatures of fill-ins of $M$ for a certain family of metrics on $M$. The proof is based on arguments in [SWW22].
@@
Let $s\in\mathbb{R}$ such that $\mathrm{scal}_{g}>s$. We define $\mathscr{R}_{g_{M},s}(M)$ as the set of Riemannian metrics $g$ on $M$ with the following property: There exists a smooth path $\zeta_{g}\colon[0,1]\to C^{\infty}({\rm Sym}(T^{*}M\otimes T^{*}M))$ of Riemannian metrics starting at $g$ and ending at $g_{M}$ such that, for all $t\in[0,1]$, we have
@@
    $\zeta^{\prime}_{g}(t)>0$ pointwise on $M$, in the sense of symmetric $(0,2)$-tensor fields,
@@
    $\mathrm{scal}_{\zeta_{g}(t)}>s$ pointwise on $M$.
@@
For any $\sigma,s\in\mathbb{R}$ and $g\in\mathscr{R}_{g_{M},s}(M)$ the following holds: Let $(\Omega_{0},g_{\Omega_{0}})$ be a fill-in of $(M,g)$ with $\mathrm{scal}_{g_{\Omega}}\geq\sigma$ and $H_{g_{\Omega}}\geq 0$. Then there exists a smooth Riemannian metric $g_{\Omega_{1}}$ on
$\Omega_{1}:=\Omega_{0}\bigcup_{\partial\Omega\approx M}\big(M\times[0,1]\big)$
with the following properties:
  • $g_{\Omega_{1}}=g_{\Omega_{0}}$ outside an arbitrarily small neighborhood of $M\times[0,1]\subset\Omega_{1}$,
  • $g_{\Omega_{1}}$ induces the metric $g_{M}$ on $\partial\Omega_{1}=M\times\{1\}$,
  • $H_{g_{\Omega_{1}}}>0$ along $M\times\{1\}$,
  • $\mathrm{scal}_{\Omega_{1}}\geq\min\{\sigma,s\}-1$,
  • $\int_{\partial\Omega_{0}}H_{g_{\Omega_{0}}}{\rm dvol}_{g}<\int_{\partial\Omega_{1}}H_{g_{\Omega_{1}}}{\rm dvol}_{g_{M}}$.
@@
Let $g\in\mathscr{R}_{g_{M},s}(M)$. Let $\zeta_{g}\colon[0,1]\to C^{\infty}({\rm Sym}(T^{*}M\otimes T^{*}M))$ be a smooth path from $g$ to $g_{M}$ that satisfies properties (i) and (ii). Consider the smooth metric $\overline{g}$ on $M\times[0,1]$ defined by
$\overline{g}=dt^{2}+\zeta_{g}(t).$
Let $\overline{A}_{t}$ be the second fundamental form and $\overline{H}_{t}=\operatorname{tr}_{\zeta_{g}(t)}\overline{A}_{t}$ be the mean curvature of $\Sigma_{t}:=M\times\{t\}\subset(M\times[0,1],\overline{g})$ with respect to the unit normal field $\frac{\partial}{\partial t}$. By (i), we have $\overline{A}_{t}>0$, hence $\overline{H}_{t}>0$ and $\overline{H}_{t}^{2}-\|\overline{A}_{t}\|^{2}>0$ for $t\in[0,1]$
Let $(\Omega,g_{\Omega})$ be a fill-in of $(M,g)$ with $\mathrm{scal}_{g_{\Omega}}\geq\sigma$ and $H_{g_{\Omega}}\geq 0$. Without loss of generality, we can assume that $\partial\Omega=M$. Applying [BH23, Proposition 3.8], we can increase the mean curvature $H_{g_{\Omega}}$ slightly, while replacing $\sigma$ by $\sigma-\tfrac{1}{2}$. Hence, we can assume that $H_{g_{\Omega}}>0$.
On $M\times[0,1]$, consider the Bartnik-type parabolic PDE
$\overline{H}_{t}\frac{\partial u}{\partial t}$ $=u^{2}\Delta_{\zeta_{g}(t)}u+\frac{1}{2}(s-\mathrm{scal}_{\zeta_{g}(t)})u^{3}+\frac{1}{2}(\mathrm{scal}_{\zeta_{g}(t)}-\mathrm{scal}_{\overline{g}})u$
@@
Since, by (ii), we have $s-\mathrm{scal}_{\zeta_{g}(t)}<0$ for $t\in[0,1]$, this equation has a positive solution $u\colon M\times[0,1]\to\mathbb{R}_{>0}$. Consider the Riemannian metric $\widetilde{g}=u^{2}dt^{2}+\zeta_{g}(t)$ on $M\times[0,1]$.
Let $H_{t}=u^{-1}\overline{H}_{t}>0$ be the mean curvature on $\Sigma_{t}$ induced by $\widetilde{g}$ with respect to the unit normal field $u^{-1}\frac{\partial}{\partial t}$. Note that $H_{0}=H_{g_{\Omega}}$. By the computation in the proof of [SWW22, Theorem 4.1] and using that $\overline{H}_{t}^{2}-\|\overline{A}_{t}\|^{2}>0$, we have
$\frac{\partial}{\partial t}\int_{\Sigma_{t}}H_{t}\,\rm{dvol}_{\zeta_{g}(t)}>0.$
@@
(4.1) $\int_{\Sigma_{1}}H_{1}{\rm dvol}_{\zeta_{g}(1)}>\int_{\Sigma_{0}}H_{0}{\rm dvol}_{\zeta_{g}(0)}=\int_{\partial\Omega_{0}}H_{g_{\Omega_{0}}}{\rm dvol}_{g}.$
Furthermore, we have $\mathrm{scal}_{\widetilde{g}}=s$.
Put
$\Omega_{1}:=\Omega_{0}\bigcup_{\partial\Omega=M\times\{0\}}M\times[0,1].$
By [BH23, Theorem 4.11] and by the choice of $u(\cdot,0)$, the mean convex singularity at the gluing region $\partial\Omega_{0}=M\times\{0\}$ can be smoothened, that is, there exists a smooth Riemannian metric $g_{\Omega_{1}}$ on $\Omega_{1}$ which is equal to $\widetilde{g}$ near $M\times\{1\}$ and whose scalar curvature is bounded below by $\min\{s,\sigma\}-1$.
The manifold $(\Omega_{1},g_{\Omega_{1}})$ is a fill-in of $(M,g_{M})$. Furthermore, by (4.1), we get
$\int_{\partial\Omega_{0}}H_{g_{\Omega_{0}}}{\rm dvol}_{g}<\int_{\Sigma_{1}}H_{1}{\rm dvol}_{\zeta_{g}(1)}=\int_{\partial\Omega_{1}}H_{g_{\Omega_{1}}}{\rm dvol}_{g_{M}}.\qed$
5. Surgery bending in Euclidean space
In this section we construct smooth hypersurfaces $\overline{\Sigma}_{\varrho}$ of size $\varrho$ in Euclidean space which we will later use to define surgery bendings with quantitative control. Let $(x_{1},\ldots,x_{k})$ be the standard coordinates on $\mathbb{R}^{k}$ and let $t$ be the standard coordinate on $[0,1]$. Equip $B_{1}^{k}(0)\times[0,1]\subset\mathbb{R}^{\,k+1}$ with the flat metric $g_{\mathbb{R}^{k+1}}=\mathrm{d}x_{1}^{2}+\dots+\mathrm{d}x_{k}^{2}+\mathrm{d}t^{2}$. Let $r:=\sqrt{x_{1}^{2}+\ldots+x_{k}^{2}}\colon\mathbb{R}^{k}\to[0,\infty)$ denote the radial coordinate.
Construction 5.1.
(Bending hypersurface) Let $\varrho\in(0,\frac{1}{10}]$ be fixed and consider the function
from Lemma A.1. Put
$a_{\varrho}\coloneqq-\xi_{\varrho}(2\sqrt{\varrho})>0.$
Fix a cutoff-function $\eta\colon\mathbb{R}\to[0,1]$ such that $\eta\equiv 1$ on $(-\infty,0]$, $\eta\equiv 0$ on $1,\infty)$, and $2\leq\eta^{\prime}\leq 0$, $|\eta^{\prime\prime}|\leq 10$. Define $\eta_{\varrho}(r)=\eta\left(\tfrac{r-\sqrt{\varrho}}{\sqrt{\varrho}}\right)$ and set
see Fig. 9. Let $G_{\varrho}\subset\mathbb{R}^{k}\times\mathbb{R}$ be the graph of the function
$B_{1}^{k}(0)\setminus B_{\varrho^{4}}^{k}(0)\to\mathbb{R},\qquad(x_{1},\dots,x_{k})\mapsto f_{\varrho}(r).$
@@ 5. Surgery bending in Euclidean space
Figure 9. The function ${f}_{\varrho}$ constructed as the interpolation of $\xi_{\rho}$ and the constant function using the cutoff-function $\eta_{\varrho}$.
Finally, put
$\overline{\Sigma}_{\varrho}:=G_{\varrho}\;\bigcup\;\{r=\varrho^{4}\}\times[a_{\varrho},1]\subset B^{k}_{1}(\varrho)\times[0,1].$
Since the inverse of the function $f_{\varrho}|_{[\varrho^{4},\varrho^{2}]}$ is the smooth function $f_{\varrho}(\varrho^{2}),a_{\varrho}]\ni t\mapsto\xi_{\rho}^{-1}(t-a_{\varrho})$ with vanishing derivative at $t=a_{\varrho}$ (see Lemma A.1), $\overline{\Sigma}_{\varrho}$ is a $C^{1}$-hypersurface that is smooth everywhere except $\{t=a_{\varrho}\}$. This $C^{1}$-singularity will be smoothened out at the end of this section.
@@
Figure 10. The hypersurface $\overline{\Sigma}_{\varrho}$ as the graph of a spherically symmetric function with a cylinder attached.
@@
Let us motivate the choice of the function $\xi_{\varrho}$. The graph of the functions $\ell_{\varrho}(r)=-2\sqrt{2\varrho}\sqrt{r-2\varrho}$ gives rise to spherically symmetric surfaces satisfying
$2\lambda_{1}=\lambda_{2}=\dots=\lambda_{k}$
for their principal curvatures, cf. Lemma A.3. In $n=3$ these are the (spatial) Schwarzschild metrics of mass $m=\varrho$. In this case the scalar curvature vanishes which does not allow us to absorb the error terms appearing in the handle construction. Therefore, we solve the ODE
$4\lambda_{1}=\lambda_{2}=\dots=\lambda_{k}$
@@
Let $\overline{\nu}_{\varrho}$ be the unit normal vector field of $\overline{\Sigma}_{\varrho}$ and denote by $\overline{\nu}_{\varrho}^{\intercal}$ the projection of $\overline{\nu}_{\varrho}$ onto the tangent space of the slices $B_{1}^{k}(0)\times\{t\}$, i.e.,
@@
The hypersurfaces $\overline{\Sigma}_{\varrho}$ from 5.1 satisfy the following properties.
@@
    $\overline{\Sigma}_{\varrho}$ is $\mathrm{O}(k)$-invariant and satisfies
    $\overline{\Sigma}_{\varrho}\cap(B_{1}^{k}(0)\times[2\varrho,1])$ $=\partial B_{\varrho^{4}}^{k}(0)\times[2\varrho,1],$
    $\overline{\Sigma}_{\varrho}\cap\Bigl(\bigl(B_{1}^{k}(0)\setminus B_{2\sqrt{\varrho}}^{k}(0)\bigr)\times[0,1]\Bigr)$ $=\bigl(B_{1}^{k}(0)\setminus B_{2\sqrt{\varrho}}^{k}(0)\bigr)\times\{0\}.$
@@
    For $r\in[\varrho^{4},\sqrt{\varrho}]$, we have
    $H(\overline{\Sigma}_{\varrho})>\frac{1}{2}|A(\overline{\Sigma}_{\varrho})|\quad\text{if }k\geq 2$
    $\mathrm{scal}(\overline{\Sigma}_{\varrho})>\frac{1}{3}|A(\overline{\Sigma}_{\varrho})|^{2}\quad\text{if }k\geq 3.$
@@
    For $r\in[\varrho^{4},\varrho^{2}]$, $k\geq 3$, we have
@@
For the final two properties let $\overline{C}\coloneqq 18k+{432}$.